# KAT-Coder Technical Report

### v2

@@ Abstract
Recent advances in large language models (LLMs) have enabled progress in agentic coding, where models autonomously reason, plan, and act within interactive software development workflows. However, bridging the gap between static text-based training and dynamic real-world agentic execution remains a core challenge. In this technical report, we present KAT-Coder, a large-scale agentic code model trained through a multi-stage curriculum encompassing Mid-Term Training, Supervised Fine-Tuning (SFT), Reinforcement Fine-Tuning (RFT), and Reinforcement-to-Deployment Adaptation. The Mid-Term stage enhances reasoning, planning, and reflection capabilities through a corpus of real software engineering data and synthetic agentic interactions. The SFT stage constructs a million-sample dataset balancing twenty programming languages, ten development contexts, and ten task archetypes. The RFT stage introduces a novel multi-ground-truth reward formulation for stable and sample-efficient policy optimization. Finally, the Reinforcement-to-Deployment phase adapts the model to production-grade IDE environments using Error-Masked SFT and Tree-Structured Trajectory Training. In summary, these stages enable KAT-Coder to achieve robust tool-use reliability, instruction alignment, and long-context reasoning, forming a deployable foundation for real-world intelligent coding agents. Our KAT series 32B model, KAT-Dev, has been open-sourced on [Uncaptioned image] https://huggingface.co/Kwaipilot/KAT-Dev.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have catalyzed a shift from static text generation toward agentic intelligence, where models autonomously reason, plan, and act within dynamic environments. In software engineering, this transformation manifests as agentic coding—a paradigm in which models function as collaborative problem solvers rather than passive code generators. Despite rapid progress, a fundamental challenge persists: bridging the gap between static, text-based training and interactive, real-world execution. Conventional code models, typically trained on massive but inert text corpora, lack the adaptive reasoning and contextual control required to operate reliably in live integrated development environments (IDEs) [1, 2, 3, 4, 5, 6, 7, 8].
Early frameworks such as Codex, CodeLlama, and DeepSeekCoder established the foundation for code generation, yet they remain limited to single-turn, instruction-following behavior. More recent efforts, including SWE-Agent[1], OpenHands[7], and Claude Code[6], introduce planning and tool-use capabilities, signaling a broader trend toward agentic execution. However, these models are often constrained by narrow domain coverage, short reasoning horizons, and homogeneous datasets that inadequately represent real software engineering workflows. As a result, their performance deteriorates when transferred from benchmark environments to production-grade systems characterized by heterogeneous tools, long-term dependencies, and frequent context shifts.
To address these limitations, we introduce KAT-Coder designed to unify reasoning, planning, and deployment robustness within a single training framework. Specifically, KAT-Coder is developed through a four-stage hierarchical curriculum that progressively enhances the model’s cognitive and operational competence:
  • Mid-Term Training — Broadens reasoning, planning, and reflection abilities through a combination of real software engineering corpora and synthetic agentic trajectories, forming a bridge between general pretraining and code-oriented supervision.
  • Supervised Fine-Tuning (SFT) — Constructs a million-sample dataset spanning over 20 programming languages, 10 development contexts, and 10 task archetypes, ensuring balanced coverage and cross-domain generalization.
@@
  • Reinforcement Learning (RL) — Employs Error-Masked SFT and Tree-Structured Trajectory Training (TST) to adapt the model to production environments with heterogeneous toolchains and non-linear context boundaries.
This curriculum reflects a closed-loop design philosophy: cognitive enrichment precedes structured supervision, which in turn grounds reinforcement learning and real-world adaptation. Through this progressive alignment, KAT-Coder evolves from a general LLM into a deployable agentic developer capable of reasoning about tasks, managing tools, and collaborating in complex software workflows.
@@ 4.1 Trie Packed Training
In this scenario, we realized that the computation of prefixes shared across multiple trajectories could be performed just once, which would significantly improve training throughput. This idea is similar to the prefix caching used for LLM inference. However, due to the involvement of back propagation during training, we cannot directly reuse cached results; doing so would ignore the gradient contributions from suffix tokens to the prefix tokens, leading to incorrect computations. Our goal is computing each shared prefix only once during both forward and backward and thus improve agentic LLM training throughput. To achieve this, we took the following steps:

### v3

@@ Abstract
Recent advances in large language models (LLMs) have enabled rapid progress in agentic coding, where models autonomously reason, plan, and act within interactive software development workflows. Yet, bridging the gap between static text-based training and dynamic real-world execution remains a core challenge. In this technical report, we present KAT-Coder, a large-scale agentic code model trained through a four-stage curriculum: Mid-Term Training, Supervised Fine-Tuning (SFT), Reinforcement Fine-Tuning (RFT), and Agentic Reinforcement Learning (Agentic RL). The Mid-Term stage enriches reasoning, planning, and reflection abilities via a corpus of real software-engineering data and synthetic agentic trajectories. The SFT stage constructs a million-sample dataset balanced across 20+ programming languages, 10 development contexts, and 10 task archetypes, enabling cross-domain generalization. The RFT stage introduces a multi-ground-truth reward formulation for stable and sample-efficient policy optimization. Building upon this, the final Agentic RL phase integrates Trie-Packed Training with Difficulty- and Entropy-Aware Advantage Rescaling, achieving efficient multi-trajectory optimization, enhanced exploration, and robust policy diversity. Together, these stages establish a unified framework that bridges reasoning enrichment, structured supervision, reinforcement alignment, and deployment robustness. KAT-Coder attains strong and balanced results across reasoning, coding, and agentic benchmarks, forming a deployable foundation for real-world intelligent coding agents. Our KAT series 32B model, KAT-Dev, has been open-sourced on [Uncaptioned image] https://huggingface.co/Kwaipilot/KAT-Dev.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have catalyzed a shift from static text generation toward agentic intelligence, where models autonomously reason, plan, and act within dynamic environments. In software engineering, this transformation manifests as agentic coding—a paradigm in which models evolve from passive code generators into collaborative problem solvers. Despite remarkable progress, a core challenge persists: bridging the gap between static, text-based training and interactive, real-world execution. Conventional code models, typically trained on massive but inert text corpora, lack the adaptive reasoning and contextual control required to operate reliably in live integrated development environments (IDEs) [1, 2, 3, 4, 5, 6, 7, 8].
Early frameworks such as Codex [3], CodeLlama [9], and DeepSeekCoder [10] established the foundation for code generation, yet they remain limited to single-turn, instruction-following behavior. More recent efforts, including SWE-Agent [1], OpenHands [7], and Claude Code [6], introduce planning and tool-use capabilities, signaling a broader trend toward agentic execution. However, these systems are often constrained by narrow domain coverage, short reasoning horizons, and homogeneous datasets that fail to capture the diversity of real software engineering workflows. Consequently, their performance deteriorates when transferred from benchmark environments to production-grade systems characterized by heterogeneous tools, long-term dependencies, and frequent context shifts.
To address these limitations, we introduce KAT-Coder—a large-scale agentic code model designed to unify reasoning, planning, reinforcement, and deployment robustness within a single training framework. Specifically, KAT-Coder is developed through a four-stage hierarchical curriculum that progressively enhances the model’s cognitive and operational competence:
  • Mid-Term Training — Broadens reasoning, planning, and reflective thinking abilities through a combination of real software engineering corpora and synthetic agentic trajectories, forming a bridge between general pretraining and code-oriented supervision.
  • Supervised Fine-Tuning (SFT) — Constructs a million-sample dataset spanning over 20 programming languages, 10 development contexts, and 10 task archetypes, ensuring balanced coverage and strong cross-domain generalization.
@@
  • Agentic Reinforcement Learning (Agentic RL) — Implements Trie-Packed Training for high-throughput multi-trajectory optimization and Difficulty- and Entropy-Aware Advantage Rescaling to maintain exploration diversity, prevent entropy collapse, and enhance policy robustness.
This curriculum reflects a closed-loop design philosophy: cognitive enrichment precedes structured supervision, which in turn grounds reinforcement alignment and culminates in real-world adaptation. Through this progressive alignment, KAT-Coder evolves from a general LLM into a deployable agentic developer capable of long-context reasoning, dynamic tool management, and collaborative problem solving across complex software workflows.
@@ 4.1 Trie Packed Training
In this scenario, we realized that the computation of prefixes shared across multiple trajectories could be performed just once, which would significantly improve training throughput. This idea is similar to the prefix caching used for LLM inference. However, due to the involvement of back propagation during training, we cannot directly reuse cached results; doing so would ignore the gradient contributions from suffix tokens to the prefix tokens, leading to incorrect computations. Our goal is computing each shared prefix only once during both forward and backward and thus improve agentic LLM training throughput. To achieve this, we proposed Trie Packed Training paradigm[30]: